BETA COEFFICIENT

Given the following information, determine the beta coefficient for Stock L that is consistent with equilibrium: = 9%; rRF = 3.9%; rM = 11.5%.

Round your answer to two decimal places.

b)

CAPM AND PORTFOLIO RETURN

You have been managing a $5 million portfolio that has a beta of 0.75 and a required rate of return of 12%. The current risk-free rate is 3.00%. Assume that you receive another $500,000. If you invest the money in a stock with a beta of 1.30, what will be the required return on your $5.5 million portfolio? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

a. Beta Coefficient

As per the CAPM, Rs = Rf + beta*(Rm - Rf)

9% = 3.9% + beta*(11.5%-3.5%)

9 = 3.9 + beta*8

beta = (9-3.9)/8 = 0.6375

Beta = 0.6375

b. First, we calculate the market return from the given information as per CAPM equation Rs = rf + beta*(Rm-Rf)

12 = 3 +0.75*(Rm-3)

12 = 3 + 0.75Rm - 2.25

Rm = 15%

Now we estimate the return of the second stock which is = Rf + beta*(Rm-rf) = 3 + 1.3*(15-3) = 18.6%

Return of stock A (Ra) = 12% and that of stock B (Rb) = 18.6%

Weight of A = 5,000,000/5,500,000 = 0.9091

Weight of B = 1-0.9091 = 0.0909

Required return on 5.5 Million = 0.9091*12 + 0.0909*18.6 = 12.60%

Required return on your $5.5 million portfolio = 12.60%